OpenStudy (keana):

subtract & simplify 5x+6/x^2-1 minus 1/x^2-1. a: x-1/x b: x+1/5 c: 5/x+1 d: 5/x-1

jimthompson5910 (jim_thompson5910):

The denominators are equal, so you can subtract the numerators and place the result over the common denominator x^2-1

OpenStudy (anonymous):

\[\frac{ 5x+6 }{ x^2 - 1 } - \frac{ 1 }{ x^2 - 1 } = \frac{ 5x + 5 }{ x^2 -1 } = \frac{ 5(x+1) }{ (x+1)(x-1) } = \frac{ 5 }{ (x-1) }\] Do what Jim said and subtract. Then notice that the denominator is a difference of two squares, so can factorise it to (x+1)(x-1) and cancel with the denominators.

OpenStudy (anonymous):

in the last sentence I meant cancel with the (x+1) on the numerator